OpenStudy (anonymous):

During a hike, Vijay saw a total of 15 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 48 wings. Use a model to determine how many birds he saw. 5 6 7 8

OpenStudy (calculusxy):

Use an equation. To do this problem, you need to do the inverse of addition because you are only trying to find how many birds are there indtead of the amount of butterflies.

OpenStudy (anonymous):

x is birds y is butterflies well make the equation like this x + y = 15 2x + 4y = 48 then we can use the substitution method y = -x + 15 2x + 4(-x + 15) = 48 2x - 4x + 60 = 48 -2x = -12 x = 6 y = 9

OpenStudy (anonymous):

so its 8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what did I say x was?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes that is the amount of birds

OpenStudy (zzr0ck3r):

what did he define x to be?

OpenStudy (anonymous):

so the answer is B

OpenStudy (anonymous):

oooo thxxx

OpenStudy (anonymous):

but diva where did you get 8??